OpenStudy (anonymous):

f(x)=2x^2-x-1. If f(x)=-1, what is x?

OpenStudy (bahrom7893):

f(x) = -1 means: -1 = 2x^2 - x - 1

OpenStudy (bahrom7893):

solve for x

OpenStudy (anonymous):

I tried to move the -1 over so that the equation would be \[2x^2-x-1+1=0\]

OpenStudy (anonymous):

Now I'm at \[2x^2-x=0\] and I'm not sure what to do next

OpenStudy (anonymous):

yes so then: -1 + 1 = 0 so: \[2x^2 - x = 0\]

OpenStudy (anonymous):

Oh, factor out an x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x(2x - 1) = 0\]

OpenStudy (anonymous):

$$f(x)=-1\\2x^2-x-1=-1\\2x^2-x=0\\x(2x-1)=0\\x=0,2x-1=0\\x\in\left\{0,\frac12\right\}$$

OpenStudy (anonymous):

Cool, those are the answers that I got. Thanks :)